PLEASE SOLVE EACH QUADRATIC EQUATION.
1. x^2+3x=-5
2. x^2-36=0
3. 2x^2-4x+10=0


Answer by edjones(8007)

2x^2-4x+10=0
x^2-2x+5=0
x^2-2x =-5
x^2-2x+1=-5+1 completing the square
(x-1)^2=-4
x-1=+-sqrt(-4)
x=1+-2i
x=1+2i, x=1-2i
.
Ed